OpenStudy (anonymous):

7.2-6). Divide as indicated. Be sure to reduce all answers to lowest terms. d^2+10d+25 d^2-25 _______________ divide _______ d+5 d-5

OpenStudy (watchmath):

\(\frac{(d+5)^2}{d+5}\cdot \frac{(d+5)(d-5)}{(d-5}=(d+5)^2\)

OpenStudy (anonymous):

r u sure this is correct.. i tried is and its wrong

OpenStudy (watchmath):

Sorry I made a mistake. It should be \[\frac{(d+5)^2}{d+5}\cdot \frac{d-5}{(d-5)(d+5)}=1\]